OpenStudy (anonymous):

Suppose that you draw two cards without replacement from a standard deck of 52 cards . Given that one card is a diamond, what is the probability that the other card is an ace?

OpenStudy (kropot72):

There are 4 aces and there are 51 cards left after drawing a diamond. P(ace) = 4/51

OpenStudy (anonymous):

book's answer is 7/51

OpenStudy (kropot72):

A standard deck of cards used for poker has only one ace in each suite, giving 4 aces. Where did the other 3 aces come from? Or, what am I missing?

OpenStudy (anonymous):

i got 4/51 as well. couldnt figure out how the book got 7/51

OpenStudy (kropot72):

Of course the card drawn first could be the ace of diamonds. In that case P(ace) = 3/51 :p

OpenStudy (anonymous):

eh

OpenStudy (anonymous):

if P(Diamond) = 13/52 (13 diamonds, 52 cards) if only draw ace, p(ace) 1/52. ~ (1/52)/(13/52) = 1/13

OpenStudy (anonymous):

?

OpenStudy (kropot72):

If the ace of diamonds is drawn first then only three aces remain for the second draw. In that case P(ace) = 3/51. If a diamond other than the ace of diamonds is drawn first then for the second draw P(ace) = 4\51 The above events are mutually exclusive, therefore the probabilities should be added giving:\[totalP(ace)=\frac{3}{51}+\frac{4}{51}=\frac{7}{51}\]

OpenStudy (anonymous):

thanks!

OpenStudy (kropot72):

You're welcome :)